OpenStudy (anonymous):

Evaluate. 102 ÷ 25 • 20 + 4 ÷ 2 – 14

OpenStudy (anonymous):

102/25= 4.08 4.08*20 = 81.6. Follow order of operations here; divide before you add. 4/2=2 81.6+2=83.6 83.6-14=69.6.